OpenStudy (anonymous):

find the local max and min using the first and second derivative test 1)f(x)= x+√1-√x 2)h(x)=lnx/√x

OpenStudy (anonymous):

The equations you posted were a bit ambiguous, so I had to guess. \[f(x) = x+\sqrt{1-x}\] \[h(x) = \frac{ \ln x }{ \sqrt{x} }\] are you able to find the respective first and second derivatives of f(x) and h(x)?